205 gram hydrogen  sulfide  reacts with 75.0 g oxygen to  produce  sulfur and  water.  Find:

a) Gram of sulfur formed

b) gram  of excess  reactant used up

c)  gram of  excess  reactant that remained  unreacted




1
Expert's answer
2021-09-09T02:32:01-0400

H2S + O2 --->>H2O + S

Molar mass of H2S = 34.1

= 205/34.1 = 6.0112 moles

Molar mass of oxygen = 31.998

75/31.998 = 2.344 moles

Mole ratio in both cases = 1:1

= 2.344+6.0112 = 8.3552 moles of Sulfur

Molar mass of sulfur = 32.065

= 32.065 × 8.3552 = 267.91g of sulfur